A cumulative product of a set (a,b,c,....) is the sequence, a, an, abc,... what is the mean of the terms in the cumulative product if (1,2,3,0)
A)0
B)2
C)2.25
D)3.5

Answer by greenestamps(13200)



If you are serious about wanting help with your question, please be polite and make sure you have presented the problem clearly.

(1) There is no "n" in the problem; clearly the sequence you mean to define is a, ab, abc,... and not a, an, abc,.... It shouldn't be our job to determine that.
(2) "...mean of the terms in the cumulative product if (1,2,3,0)." That is not English. We can guess that the meaning is "... cumulative product OF (the set) (1,2,3,0)". But again, we shouldn't have to be guessing that; you should make it clear in your statement of the problem.

And now that we have gotten past the sloppy presentation of the problem, what about the problem is giving you difficulty?

Using the definition you give, the cumulative product is
a = 1
ab = 1*2 = 2
abc = 1*2*3 = 6
abcd = 1*2*3*0 = 0

The mean of the terms is the sum of the four terms, divided by the number of terms, which is 4. That calculation gives one of the given answer choices.